Calculate the number of moles, molecules atoms, bonds , electrons , protons , neutrons, in lone pair present in 1 litre water (Density of water = 1ml)

Grade:11

1 Answers

Umashankar Kumar
6 years ago
Density of water=1g/ml ,volume=1lit ;mass=1kg. ,No of moles=1000/18=55.56. no of molecules=moles×avogadro no=55.56×6.023×10to power 23=3.36×10topower25. No of atoms=no of molecules×no of atoms present in water=1.008×10to power26. No of bonds=no of molecules×no of bonds in one molecule of water=3.36×10to power25×2=6.72×10to power25. No of electrons=no of electrons in one molecule of water×no of molecules=3.36×10to power26=no of protons. No of neutrons=2.688×10to power26. No of LP of electrons=6.72×10to power25.